# Self-Specialization: Uncovering Latent Expertise within Large Language Models

### v1

@@ 4.2 Results
In Table 1, we present the comparative results of our self-specialized MPT-30B model against its base counterpart across 10 distinct biomedical NLP tasks. The evaluation is conducted using various k-shot prompting to analyze the impact of different numbers of in-context examples on model performance with/without specialization.
@@
Our findings reveal that the self-specialized MPT-30B model exhibits remarkable progress in the majority of tasks across all configurations, yielding a surprisingly substantial (up to 18 points) improvement in average scores. Specifically, the scores () witness a rise from 25.15 to 36.63 in a zero-shot setting, from 26.65 to 42.29 in a 1-shot setting, and from 30.18 to 48.41 in a 5-shot setting, respectively. Importantly, the effectiveness of self-specialization becomes evident as it uncovers the latent expertise encoded within the “generalist” base model, showcasing the potential of leveraging inherent knowledge for enhanced domain-specific performance. These advancements underscore the self-specialized model’s versatility and adaptability in addressing a wide array of tasks present in a specialized domain.
@@
How does it compare against larger/generally aligned models?
@@
In Figure 4, we compare our self-specialized MPT-30B model with 65B models, including LLaMA-65B, and its general instructions aligned variants. Interestingly, the results reveal that our model, despite its 2.2x smaller size, surpasses all 65B models. This not only highlights the lower expert domain performance trade-offs of the “generalist” models in terms of encoding a vast array of general knowledge into a finite set of parameters but also underscores the effectiveness of our (data and parameter efficient) approach to model specialization. Moreover, the efficiency and practicality of our simple self-specialization are further reinforced by the fact that the model is trained using only 5K22252K for Alpaca and 360K for Dromedary instruction data self-produced with minimal (only 80333175 for Alpaca and 195 for Dromedary) seeds. This training process, facilitated by the incorporation of QLoRA, which adds only 0.14% trainable parameters to an otherwise frozen model, only takes a few hours on a single (A100 80GB) GPU.
Model -Score Rouge-L 2nd Iter. Self-Specialized MPT-30B 36.63 34.79 w/ Same Instruction Set 35.82 34.20 1st Iter. Top-5 Docs 34.57 32.88 Top-1 Docs 29.65 27.90 No Docs 33.72 32.14 Base MPT-30B 25.15 23.75
@@ 4.3 Ablations & Analyses
We investigate the influence of incorporating domain-specific corpus like PubMed in the response generation phase, which enriches the model with pertinent biomedical information. As observed in the “1st Iter.” section of Table 2, there is a notable variation in performance depending on the number of documents incorporated. Our findings indicate that the use of the top-5 documents yields the best results. Interestingly, incorporating only the top-1 document appears to degrade the performance, a phenomenon we conjecture is due to the noise originating from an imperfect retriever. Conversely, employing top-5 documents with probability marginalization (eq. 2) seems to mitigate this issue, enabling the model to exploit informative knowledge.
@@
In Section 3.5, we discussed the potential of employing an iterative process by leveraging the self-specialized model instead of the base model throughout the generation process. As evidenced in Table 2, initiating a “2nd Iter.” of self-specialization results in further performance enhancement. Additionally, we consider two scenarios differentiated by whether the same instruction set is used to train the self-specialized model in the first and the second iterations. Our findings show that employing a distinct set of instructions for the second iteration in response generation is more effective. This could potentially be attributed to the limitation of using a confined generated instruction set, which might limit the model’s generalization capabilities.
Can self-specialization also be applied to a different model (or model size)?
To demonstrate the applicability of self-specialization other than to the MPT model, we apply it to another open-source model, Falcon-40B. Figure 5 shows the results on five different datasets. Notably, we observe a trend of improvement analogous to that of MPT-30B when self-specialization was applied to Falcon-40B, thereby substantiating that the technique is not exclusive to the MPT model. Surprisingly, despite its larger parameter size, the Self-Specialized Falcon-40B underperforms its MPT-30B counterpart, while at the same time significantly improving upon the base Falcon-40B.
@@
To quantitatively assess the quality of the data generated through self-specialization, we train a model using 3.7K instances of available human-labeled data in a multi-task learning setting and compare its performance to that of a model trained on 5K instances of generated synthetic self-specialization data, as depicted in Figure 5. Although the model trained on supervised data exhibits higher performance as expected, the performance gap between the two models is not large, further underscoring the effectiveness of the proposed self-specialization. In Figure 6, we showcase a qualitative visualization that analyzes the synthetic data generated through self-specialization. Additionally, some examples are provided in Table 4 & 5 in Appendix, offering insights into the quality of the self-generated specialization data.
@@ 5 Related Work
The goal of instruction-tuning and alignment of large language models (LLMs) is to achieve cross-task generalization or to align with human preferences. This can be accomplished by either training LLMs directly with human-labeled data (Ouyang et al., 2022; Wei et al., 2022; Mishra et al., 2022; Wang et al., 2022b) or data generated by larger models (i.e., distillation) (Taori et al., 2023; Chiang et al., 2023). Recent studies have shown that LLMs are self-instructors. Wang et al. (2022a) showed that with in-context prompts, GPT-3 (Brown et al., 2020) can generate high-quality instruction-responses pairs for its own alignment. Sun et al. (2023) further suggests that using principles can minimize human supervision while covering a broad spectrum of scenarios with the open-source model, LLaMA-65B (Touvron et al., 2023). While enhancing general alignment, according to our presented evidence, these approaches are unlikely to induce specialization in expert domains, leaving different domain expertise in “superposition” inside the model. To the best of our knowledge, we are the first to show the potential techniques for expert domain specialization through self-alignment, effectively “uncovering” a domain expert out of the model in a parameter- and data-efficient manner.
@@
6 Discussion
While our study provides encouraging insights into the capabilities of self-specialization, this is an initial step in opening up new opportunities. We recognize that there is much to learn and explore in this exciting direction as discussed below. The promising results, achieved even with the proposed simple scheme, suggest that further refinement of this approach and exploration across diverse specialized domains could be pivotal, contributing to the ongoing efforts to uncover the embedded expertise of (generalist) large language models. In what follows, we discuss a few noteworthy considerations and potential directions.
On the scale of a base model and data.
The study focuses on employing a large base model size (i.e., 30B) for the initial exploration, motivated by the success of preceding research (Wang et al., 2022a; Sun et al., 2023) that employed general self-alignment with even larger models (e.g., 65B, 175B). Nonetheless, we believe exploring the potential of smaller-scale models (e.g., 7B) as a base for self-specialization presents an additional intriguing avenue for future research, possibly enhancing the practical feasibility. Regarding the quantity of synthetic data generated through self-specialization, we constrained it to 5K for the sake of simplicity and efficiency (Zhou et al., 2023). While this led us to highlight the data efficiency of our model, future studies could investigate the extent to which increasing the data could further enhance the expertise.
Additional domains.
While our study primarily focuses on the biomedical domain, the applicability and effectiveness of self-specialization in other specialized domains, such as sports, remain an open avenue for exploration. As an initial effort, we present a case study of a self-specialized model on sports in Table 6 & 7, along with the visualization of generated data in Figure 7. We hope that this could offer insights into the versatility of self-specialization, although the model is not yet perfect, and thorough evaluations are required in future work. Different domains inherently pose unique requirements and nuances, and understanding how self-specialization adapts to these variations is a valuable direction for future work.
Mixture-of-self-aligned-experts.
The potential demonstrated by self-specialization in our study paves the way for another fascinating research direction: the combination of distinct self-specialized models. This would involve integrating the expertise of various self-specialized models, aiming to create a more comprehensive and adaptable solution. Investigating the synergies and challenges in combining different specialized models (e.g., lightweight specialized LoRA layers) can lead to the development of a model with enhanced concurrent proficiency across a broader spectrum of specialized domains (at once), offering a holistic approach to maximizing the extraction and utilization of the latent expertise of LLMs.
@@ 7 Conclusion
Our exploration into self-specialization, drawing inspiration from the recent achievements in general-purpose self-alignment (Wang et al., 2022a; Sun et al., 2023), aimed to elucidate the latent expertise within large language models (LLMs) with very limited human supervision. This scheme, which incorporates a few domain-specific seeds and external knowledge into the synthetic data generation process, demonstrated promising results in a specialized domain. The self-specialized model (30B) exhibited remarkable performance, outshining its base model, MPT-30B, and even surpassing larger existing generally aligned models (65B). This illuminates the intrinsic challenges of encoding vast general knowledge into limited parameters and underscores the efficiency of self-specialization. Remarkably, the model’s efficient training, marked by minimal data usage and the integration of QLoRA (Dettmers et al., 2023), adds another layer to its practicality in terms of parameter and data efficiency. These findings signify a promising advancement in the field, suggesting a pathway for leveraging inherent domain-specific expertise in LLMs and offering a large variety of exciting opportunities for future work in self-specialization.
@@ Appendix A Explanations of Evaluation Datasets
Below are brief descriptions for each dataset.
@@
This is a biomedical QA dataset that necessitates models to produce answers from given questions and corresponding contexts within the biomedical domain. There are three distinct subsets that can be divided according to question types: Factoid, List, and Yesno.
@@
PubMedQA is another biomedical QA dataset featuring research questions along with their corresponding abstracts and answers sourced from PubMed444https://www.ncbi.nlm.nih.gov/pubmed. To diversify the task types, we focus on a long-form answer (i.e., conclusion).
AnatEM (Pyysalo & Ananiadou, 2013).
This is a Named Entity Recognition (NER) task for anatomical entities in biomedical texts. Models are tasked with identifying all anatomy-named entities and their corresponding types from given a small paragraph.
@@
The Cancer Genetics (CG) is an information extraction task targeting the recognition of events in text, encompassing multiple levels of biological organization, from molecular to whole organisms.
@@
The NCBI disease corpus, derived from the National Center for Biotechnology Information, focuses on disease name recognition.
@@
The Drug-Drug Interaction (DDI) dataset is tailored for identifying interactions between different drugs in biomedical texts. Following Parmar et al. (2022), this work considers only binary Relation Extraction (RE), determining whether there is an effect of given two drugs.
@@
This is a Sentiment Analysis (SA) dataset that is required to predict the sentiment of individuals towards medical drugs. Specifically, given a text and a drug, a model determines the effect of the drug as “positive”, “negative”, or “neutral”.
@@
The Hallmarks of Cancer (HoC) dataset is curated for classifying (zero to many) biomedical texts related to cancer into categories representing different hallmarks of cancer. In particular, these hallmarks include “sustaining proliferative signaling”, “resisting cell death”, “genomic instability and mutation”, “activating invasion and metastasis”, “tumor promoting inflammation”, “evading growth suppressors”, “inducing angiogenesis”, “enabling replicative immortality”, “avoiding immune destruction” and “cellular energetics”.

### v2

@@ 5.1 Comparison with Baselines
In Table 2, we present the comparative results of our self-specialized model against its base counterpart across 10 distinct biomedical NLP and 10 financial NLP datasets. The evaluation is conducted with varying numbers of in-context demonstrations, k.
@@
Our findings reveal that the self-specialized model exhibits remarkable progress in the majority of tasks across all configurations in both domains, yielding a substantial (up to 18 points) improvement in average scores. Specifically, the average scores () in biomedicine rise from 30.18 to 48.41 in a 5-shot setting.222Even excluding BioASQ-Yesno as an outlier due to the base model’s low performance, self-specialization still shows significant gain over the base model: 32.55 to 43.21 (5-shot). Appendix C.3 includes the detailed discussion. In finance, the improvements are 14.53 (0-shot), 6.77 (1-shot), and 12.86 (5-shot), respectively. These advancements in both domains underscore the self-specialization’s generalizability in addressing a wide array of tasks across different specialized domains.
Imact on ICL capability.
A potential concern on self-specialization tuning is its impact on the base LLM’s in-context learning capabilities, as we did not tune the model with demonstrations. Comparing the capabilities before and after self-specialization, the improvement after adding demonstrations (from 0 to 5) of our self-specialized model on biomedicine in Table 2 is 36.63 to 48.41 (=11.78), while that of the base model is 25.15 to 30.18 (=5.03), indicating even better ICL capability with in-domain knowledge acquisition.
Performance drop on some tasks.
Our analysis does identify a few instances where performance drops as shown in Table 2. This indicates room for further refinement, especially for tasks like ConvFinQA that require a set of specific capabilities beyond mere domain knowledge. We evidenced that a minor proportion ( 2%) of generated data partially resembles ConvFinQA, due to our generation’s nature involving creative brainstorming for diversity. The specific demands of ConvFinQA, including numerical reasoning, structured tables, and conversations extend beyond basic domain knowledge and were insufficiently covered within our dataset. This gap likely contributes to the observed performance trade-offs.
However, we re-emphasize that there are significantly bigger gains in many of the cases (e.g., 45 out of 60 experiments across datasets and k), outweighing the regression overall. Acknowledging the inherent variability of in-context learning (Min et al., 2022), we present the variances with 5 different sets of demonstrations in Figure 4 based on LLaMA-2-7B in biomedicine, showing significant average improvements of 8.25 () and of 14.42 ().
@@
How does self-specialization compare against larger/generally aligned baselines?
@@
In Figure 3, we compare our self-specialized MPT-30B model with 65B models, including LLaMA-65B, and its general instruction aligned variants (e.g., Alpaca based on Self-Instruct) in the biomedical domain. Interestingly, the results reveal that our model, without extensive data, surpasses all baselines, including 65B models, despite its 2.2x smaller size. This not only highlights the lower expert domain performance trade-offs of the “generalist” models in terms of encoding vast general knowledge into a finite set of parameters, but also underscores the effectiveness of our parameter-efficient approach to model specialization. We also show that our self-specialized model outperforms the supervised general-purpose model, MPT-30B-instruct in all tasks, which highlights the benefits of in-domain instruction data. Moreover, as a reference point, we present a comparison with a fully-supervised SOTA model that is fine-tuned on the biomedical datasets in Table 7, contextualizing our progress to better understand practical utility, discussed in Appendix C.4. Notably, the data efficiency of our simple self-specialization is further reinforced by the fact that the model is trained using only 5K33352K for Alpaca and 360K for Dromedary. instruction data self-produced with minimal (only 80) seeds.444175 for Alpaca and 195 for Dromedary. This training process, facilitated by the incorporation of QLoRA, adding only 0.28% trainable parameters to an otherwise frozen model, only takes a few hours on a single GPU (A100 80GB).
Model -Score Rouge-L w/ Top-5 Docs 34.57 32.88 w/ Top-1 Docs 29.65 27.90 w/o Retrieval 33.72 32.14 Base MPT-30B 25.15 23.75
@@ 5.2 Ablations & Analyses
We investigate the influence of incorporating a domain-specific corpus like PubMed in the response generation phase. Table 3 shows optimal results with the top-5 documents, while using just the top-1 document decreases performance, likely due to noise from an imperfect retrieval process, aligned with findings from previous work (Yoran et al., 2023) that adding irrelevant (i.e., random) context dramatically decreases performances. Conversely, employing the top-5 documents with probability marginalization (eq. 3.3) seems to mitigate this issue, enabling the model to exploit informative knowledge. Interestingly, we observe that self-specialization demonstrates strong performance even without retrieval, suggesting domain-relevant knowledge is intermingled with other information acquired during pre-training, which self-specialization uncovers to better utilize. Given this, the added complexity of retrieval mechanisms, though potentially advantageous, emerges as optional within our framework.
Model -Score Rouge-L w/ Iterative Process 36.63 34.79 Self-Specialization 34.57 32.88 Base MPT-30B 25.15 23.75
@@
In Section 3.5, we discussed the potential of employing an iterative process by leveraging the self-specialized model instead of the base model throughout the generation process. Table 4 shows the ablation study, where each iteration involved generating 5K samples, and final results were obtained using 5K samples from the last iteration for a fair comparison. We observe that the iterative process leads to further performance enhancements, compared to the one w/o iteration. In our preliminary tests, we rarely find meaningful improvements with the subsequent iteration, which we leave for future work to refine.
Self-specialization vs. domain pre-training.
We compare our model based on LLaMA-2-7B with existing baselines (Wu et al., 2023): MedLLaMA-13B and PMC-LLaMA-7B/-13B. The former is an LLaMA variant further pre-trained on a large domain-specific corpus (i.e., medicine), and the latter is further instruction-tuned using annotated/synthetic datasets, including medical QA, rationale for reasoning, and conversational dialogues. Notably, we find that our self-specialized 7B model is on par with or better than MedLLaMA-13B () and PMC-LLaMA-13B () despite their larger parameters and extensive domain-specific tuning. Additionally, using our 7B-generated data to specialize MedLLaMA indicates that self-specialization can enhance domain-specific pre-training (), suggesting complementarity.
Impact of the number of self-generated data.
In Figure 5, we analyze the impact of the number of self-specialization data within biomedical and financial domains. Starting from zero, a sharp increase in score is observed as we introduce the first 100 instances which largely consist of seed instructions, underlining the significant impact of seeds not only as in-context demonstrations but also as training data. The performance continues to rise steadily with additional data, plateauing around 5K instances, supporting our decision on the use of 5K data. Self-specialization’s success with relatively small self-generated data highlights its data efficiency and practicality.
@@
In Figure 6, we showcase a qualitative visualization that analyzes the synthetic data generated through self-specialization, confirming that self-specialization produces domain-focused data. To quantitatively assess the quality, Figure 7 in Appendix compares our model against a model trained on labeled data, which shows a narrow performance gap, implying the quality of generated data. Additionally, some examples are provided in Table 11 & 12 in Appendix, offering insights into the quality of the self-generated specialization data.
@@ 6 Related Work
The goal of instruction-tuning and alignment of large language models (LLMs) is to achieve cross-task generalization or to align with human preferences. This can be accomplished by either training LLMs directly with human-labeled data (Ouyang et al., 2022; Wei et al., 2022; Mishra et al., 2022; Wang et al., 2022b) or data generated by larger models (i.e., distillation) (Taori et al., 2023; Chiang et al., 2023). Recent studies have shown that LLMs are self-instructors. Wang et al. (2022a) showed that with in-context prompts, GPT-3 (Brown et al., 2020) can generate high-quality instruction-responses pairs for its own alignment. Sun et al. (2023) further suggests that using principles can minimize human supervision while covering a broad spectrum of scenarios with the open-source model, LLaMA-65B (Touvron et al., 2023a). While enhancing general alignment, according to our presented evidence, these approaches are unlikely to induce specialization in expert domains, leaving different domain expertise in superposition inside the model. To the best of our knowledge, we are the first to show the potential for expert domain specialization through self-alignment, effectively “uncovering” a domain expert out of the model in a parameter- and data-efficient manner.
@@ 7 Conclusion
Our exploration into self-specialization aimed to elucidate the latent expertise within large language models (LLMs) with limited human supervision. This scheme demonstrated promising results in specialized domains. The self-specialized model exhibited remarkable performance, outperforming its base model, MPT-30B, and even surpassing larger generally aligned models (65B). This illuminates the intrinsic challenges of encoding vast general knowledge into limited parameters and underscores the efficiency of self-specialization. Remarkably, the model’s efficient training, marked by minimal data usage and the integration of QLoRA (Dettmers et al., 2023), adds another layer to its practicality in terms of parameter and data efficiency. These findings signify a promising pathway for leveraging inherent expertise in LLMs and offering a large variety of exciting opportunities for future work.
Limitations
While our study provides encouraging insights into the capabilities of self-specialization, this is an initial step in opening up new opportunities. We acknowledge the need for further exploration and note some limitations and considerations.
Sensitivity of in-context learning.
In Table 2, we observed that performances sometimes dropped with more in-context learning demonstrations. While recognizing, the performance fluctuation is not an issue stemming from our self-specialization tuning, as it happens for the base LLM as well as GPT-3 (Appendix C.2) , demonstrating an inherent challenge in in-context learning. This phenomenon is not unique to our self-specialization approach, but a broader challenge in the field.
Training and generation strategies.
We avoided using demonstrations during training (Min et al., 2022) to maintain flexibility in the number of examples available during inference. We aimed to ensure that zero-shot performance remains unaffected by tuning to rely on demonstrations.
Unlike previous work (Wang et al., 2023a) that generates instructions first and then inputs/responses together, our approach simultaneously generates instructions and inputs, followed by responses. This strategy, inspired by a more recent work (Sun et al., 2023), enables the use of inputs as queries for retrieval prior to response generation. Despite the specific reasons outlined above, we recognize the potential of the alternative strategies as avenues for future exploration, which can be orthogonal to our current approach.
Filtering.
In our method, we opted not to implement an automatic filtering process for the generated data. In a preliminary study to assess feasibility, we attempted to filter out low-quality data manually, however we did not observe a noticeable improvement. We hypothesized that incorporating this seemingly unuseful data may even enhance the model’s robustness by preventing overfitting to those generated data. Despite this, we acknowledge the importance of further investigating filtering techniques for potential improvements.
Potential data contamination and bias propagation.
Being cautious with potential data contamination from base language models during self-specialization, we conducted stringent measures following practices in GPT-3 (Brown et al., 2020) and PaLM Chowdhery et al. (2022). We adopted n-gram overlap analysis (with n=8 and a threshold of 70%) to scrutinize similarities between our generated data and all the test sets, revealing no significant overlaps. Moreover, a detailed manual inspection of 200 random instances corroborated this finding. When concerned about retrieval sources, one can apply the n-gram overlap filtering, though our sources are PubMed abstracts without explicit labels, which inherently ensures little risk of data overlap. Meanwhile, we acknowledge the inherent risk of propagating biases from the pre-trained data.
Acknowledgments
This research is supported in part by the NSF (IIS-2052498), ODNI and IARPA via the HIATUS program (contract 2022-22072200004). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Explanations of Evaluation Datasets
Below are brief descriptions for each dataset in biomedical and financial domains. All datasets are in English.
A.1 Biomedicine
@@
This is a biomedical QA dataset that necessitates models to produce answers from given questions and corresponding contexts within the biomedical domain. There are three distinct subsets that can be divided according to question types: Factoid, List, and Yesno. This dataset is publicly available upon a data use agreement. The data are originally intended to be used as training and development data, and we use the small part of the training set as seeds (i.e., 5 seeds), and the test set for evaluation (500 for each question type). CC BY 2.5.
@@
PubMedQA is another biomedical QA dataset featuring research questions along with their corresponding abstracts and answers sourced from PubMed555https://www.ncbi.nlm.nih.gov/pubmed. To diversify the task types, we focus on a long-form answer (i.e., conclusion). We use 5 labeled data for seeds and 500 for evaluation. MIT license.
AnatEM (Pyysalo and Ananiadou, 2013).
This is a Named Entity Recognition (NER) task for anatomical entities in biomedical texts. Models are tasked with identifying all anatomy-named entities and their corresponding types from given a small paragraph. Non-commercial purposes only. 404 test data are used for evaluation and 5 training instances are used for seeds. CC BY-SA 3.0.
@@
The Cancer Genetics (CG) is an information extraction task targeting the recognition of events in text, encompassing multiple levels of biological organization, from molecular to whole organisms. 5 training data are used for seeds, and the number of evaluation data is 200. CC BY-SA 3.0.
@@
The NCBI disease corpus, derived from the National Center for Biotechnology Information, focuses on disease name recognition. According to the annotation guideline of this dataset, organism names such as humans, and also gender are excluded for annotation. We use 5 training instances for seeds, and 100 for evaluation. The data is freely available to the public for use. CC0 1.0 license.
@@
The Drug-Drug Interaction (DDI) dataset is tailored for identifying interactions between different drugs in biomedical texts. Following Parmar et al. (2022), this work considers only binary Relation Extraction (RE), determining whether there is an effect of given two drugs. The data cannot be used for any commercial purposes. We use 5 data for seeds, and 500 for evaluation. CC BY-NC 4.0.
@@
This is a Sentiment Analysis (SA) dataset that is required to predict the sentiment of individuals towards medical drugs. Specifically, given a text and a drug, a model determines the effect of the drug as “positive", “negative", or “neutral". 5 training instances are used for the seed construction, and 500 test set for evaluation. The license is unknown.
@@
The Hallmarks of Cancer (HoC) dataset is curated for classifying (zero to many) biomedical texts related to cancer into categories representing different hallmarks of cancer. In particular, these hallmarks include “sustaining proliferative signaling", “resisting cell death", “genomic instability and mutation", “activating invasion and metastasis", “tumor promoting inflammation", “evading growth suppressors", “inducing angiogenesis", “enabling replicative immortality", “avoiding immune destruction" and “cellular energetics". The number of evaluation data is 200 and 5 training data are used for seed demonstrations. GPL-3.0 license.
A.2 Finance
EDT-Summarization (Zhou et al., 2021).
This dataset challenges models to perform abstractive summarization on financial news articles, condensing detailed information into succinct summaries. 8 training instances are used for seeds, and 500 instances for evaluation. This data is publicly available.
InsuranceQA (Feng et al., 2015).
This is an open-book question-answering task about insurance, demanding models to extract and provide specific insurance-related information. Seed demonstrations include 8 training data and the number of evaluation instances is 500. This dataset is provided as is and for research purposes only.
ConvFinQA (Chen et al., 2022).
This is a dataset for conversational question-answering over financial report tables, testing a model’s ability to reason and respond within a conversational context. We use 8 training data for the seed construction, and evaluation uses 500 test instances. MIT license.
Fin3 (Salinas Alvarado et al., 2015).
This is a financial NER dataset based on financial agreements to aid credit risk assessments. 8 training data are used for seeds and 100 test data for evaluation. CC-BY 3.0.
FiNER_139 (Loukas et al., 2022).
This NER task focuses on financial texts, where models identify and classify financial-related entities like numbers. This dataset includes a much larger label set of 139 entity types. Seed data encompass 8 training instances and the number of test data is 500. MIT license.
KPI-EDGAR (Deußer et al., 2022).
Models are tasked with extracting key performance indicators (KPIs) from financial documents. Categories for KPIs include current and previous year values, annual changes, subordinate and descriptive attributes, co-references, and false-positive. We use 212 test instances for evaluation and 8 training instances for seed demonstrations. MIT license.
EarningsCall (Roozen and Lelli, 2021).
This is a binary sentiment analysis task where models evaluate sentiments from stock values and transcripts of earnings calls, reflecting the financial sentiments expressed. 8 training instances are used for seeds, and 500 test set for evaluation. CC0 1.0 license.
Financial_Phrasebank (Malo et al., 2014).
This dataset involves (3-way) sentiment analysis of financial news headlines, assessing the underlying sentiment conveyed by the language used. Commercial uses of this data may be allowed upon contacting the authors. 8 training data and 500 test data used for seeds, and evaluation, respectively. CC BY-NC-SA 3.0.
FIQA-SA (Maia et al., 2018).
It consists of aspect-based sentiment analysis tasks within financial texts, requiring models to discern sentiment regarding specific aspects mentioned. The number of evaluation data is 234 and seed demonstrations include 8 training instances.
Gold Commodity News (Sinha and Khandait, 2021).
This dataset involves classifying financial news headlines about gold commodities into categories such as market movement direction or type of financial news (e.g., direction up, down, pastprice, futurenews, etc). The seed data includes 9 binary-class version and also 9 multi-class version of training set, and evaluation uses 500 multi-class version of test data. The license of this data indicates data files © original authors.
Biomedicine Worst Average Best Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized QA BioASQ-Factoid 30.90 37.35 43.47 50.00 51.96 57.61 BioASQ-List 35.09 42.17 42.91 44.57 47.57 46.99 BioASQ-Yesno 8.80 85.27 13.60 91.49 21.20 95.20 PubMedQA 11.98 24.16 24.19 26.78 31.69 31.31 NER AnatEM 6.59 11.99 7.93 16.33 9.63 21.25 BioNLP13CG 21.76 24.93 24.19 32.63 26.03 41.16 NCBI 17.99 14.35 21.44 34.67 27.88 46.54 RE DDI 49.20 49.40 49.86 51.47 51.00 53.40 SA Medical Drugs 11.40 32.80 19.27 51.07 35.00 65.80 DC HoC 2.44 6.01 26.40 25.42 62.84 62.65 Average 19.62 32.84 27.33 42.44 36.48 52.19
Finance Worst Average Best Task Dataset Base Self-Specialized Base Self-Specialized Base Self-Specialized SUM EDT-Summarization 6.40 21.90 11.41 23.15 13.97 24.00 QA InsuranceQA 3.03 19.87 6.51 22.67 9.96 24.36 ConvFinQA 15.74 5.25 22.07 12.66 28.77 20.88 NER Fin3 6.80 23.93 8.09 31.58 9.94 43.87 FiNER_139 10.24 14.84 30.45 25.43 44.34 35.63 RE KPI-EDGER 11.22 31.02 34.65 49.49 49.46 63.90 SA EarningsCall 46.80 47.74 48.88 48.18 50.08 48.80 Financial_Phrasebank 9.4 47.60 20.73 63.20 29.20 73.20 FIQA-SA 44.44 56.84 54.84 62.82 61.54 70.09 CLS Gold Commodity News 21.95 43.03 40.77 53.10 61.93 61.20 Average 17.60 31.20 27.84 39.23 35.99 46.59